OpenStudy (anonymous):

chip earns a base salary of 500 per month as a salesman in addition to the salary he earns 90 per products that he sells. if his goal is to earn 5000per month how man products does he need to sell

OpenStudy (anonymous):

500 + 90x = 5000 x = 4500/90 = 50 products